STORM 2-28-11

Laser powers:

  • 750: 104 mW
  • 647 : 150 mW
  • 561: 150 mW
  • 488: 150 mW
  • 405: 1.5 mW

Sample 2H.

  •  750 switching nicely.
  • 647 very bright conventional images.  Under high laser power image still inverts — low nuc singal high background.  Background does not remain high, but not sure that localizations are coming from nuclei.
  • 561 much like 647.  poor localizations

Sample rad21

  • not fully etched from 647 / 561 view.
  • lower laser power doesn’t excite the ‘resin?’ background as much and can still switch maybe.
  • Etched new slide rad21 for 3 hrs, some of sample beginning to detach.  Identically treated 2H (had more 70% EtOH rinse at the end but otherwise identical), is completely gone after 3 hrs.
O/N run of H2.  (K27ac-750, H3-647, K27me3-Cy3B [561]).   
  • 75o channel behaves as expected.  
  • Not sure about H3 and K27 in this sample.  Still worried about resin.  
  • Dyes do photoswitch in all channels, lets see how it looks tomorrow. 

new slide rad21

  • 647 keeps switching for quite a while (51K frames and counting)  We’ll see how clustered it is.  750 looked alright but didn’t switch very long. But I think I exhausted the 750 H3 secondary with repeated reuse.
  • 561 Lots of bright switching events but not sure pattern is correct.  Dot shape is more irregular.

single section of young embryo Dm01 (nice round nuclei).

  • 647 very strong for finding embryos.
  • 750 looks good but sparse number of localizations.
  • 647 going strong, good for 20K frames
  • 561 didn’t look very good by confocal in resin block.  Briefly looks like some dye is localized during initial activation.  plenty of switching.  Dots irregular size and shape though — might be background.  Higher threshold contrast looks better.
  • 488 looked strong by confocal in resin block.  Let’s see if we can turn this into a reaonsable channel.
This entry was posted in Microscopy and tagged . Bookmark the permalink.